# Alternating Perception-Reasoning for Hallucination-Resistant Video Understanding

### v1

@@ Abstract
Sufficient visual perception is the foundation of video reasoning. Nevertheless, existing Video Reasoning LLMs suffer from perception shortcuts, relying on a flawed single-step perception paradigm. This paradigm describes the video and then conducts reasoning, which runs the risk of insufficient evidence and emergent hallucinations. To address these issues, we introduce a new framework that integrates a loop-based paradigm with an anti-hallucination reward. First, to address the insufficient evidence, we introduce the Perception Loop Reasoning (PLR) paradigm. Instead of describing the video at once, each loop requires the model to describe a video segment with precise timestamps, analyze this segment, and decide the next action. Second, for the risk of hallucinations, the Factual-Aware Evaluator (FAE) evaluates each perception result as a reliable anti-hallucination reward. This reward encourages the model to provide sufficient and precise video evidence. Our FAE, which performs comparably to GPT-4o, is tuned on our AnetHallu-117K, a large-scale hallucination judgment preference dataset. Extensive experiments show that our Video-PLR achieves the state-of-the-art in both 3B and 7B parameter scales and has the best data efficiency. Our code, models, and datasets are released on: https://github.com/BoweiPu/VideoPLR.
@@ 1 Introduction
Video Large Language Models (VideoLLMs) [llava-video, liLLaVAOneVisionEasyVisual2024, wangInternVideo2ScalingFoundation2024a] demonstrate remarkable potential in traditional video understanding. Building on this success, researchers are increasingly turning their attention to complex video tasks. Inspired by the Think-then-Answer paradigm [guo2025deepseek], many VideoLLMs [wang2025videorft, zhang2025tinyllava, feng2025video] tackle complex problems by generating intermediate reasoning steps. This reasoning is based on video perception: the accurate description of visual information in the video. Thus, insufficient perception or hallucinations often lead to flawed or erroneous reasoning.
@@
Realizing this multi-step perception is challenging, as current models lack the ability to re-perceive, and no dataset exists to support it. Following the PLR paradigm, we construct VideoPLR-14K, a novel dataset designed to bootstrap the model to generate responses in the PLR format, encompassing both sequential causal reasoning and non-sequential segment ordering. Furthermore, anti-hallucination is the key to precise perception, which is the foundation of reasoning. However, previous models, including Open-source VideoLLMs [wang2024videohallucer, li2025vidhalluc] and vision-language models [zhai2023sigmoid], cannot provide reliable factual evaluation. To fill this gap, we introduce a Factual-Aware Evaluator (FAE), which is meticulously trained on our newly constructed AnetHallu-117K dataset. The ANetHallu-117K is constructed from the human-annotated ANet [caba2015activitynet] and contains 5 types of hallucination and hybrid CoT binary classification preference data, which can reduce discriminative bias. FAE is comparable to GPT-4o in evaluating text-video consistency to serve as a reliable anti-hallucination reward for RL training. The segmented approach of PLR is also more reliable for hallucination evaluation. Specifically, their constrained visual context simplifies grounding the descriptions, making anti-hallucination more readily identifiable.
These two efforts jointly constitute our final model, Video-PLR via reinforcement learning [guo2025deepseek], achieving high data efficiency. For the RL stage, we utilize only 6K of mixed data, including five types of tasks. Our models (3B and 7B) achieve SOTA results on multiple benchmarks, as illustrated in Fig. 1(c). Compared with previous models, Video-PLR integrates accurate timestamps directly into the reasoning text, enabling a unified consideration of temporal grounding. It is worth mentioning that by simply adding 0.5K of Video Temporal Grounding data, we outperform models trained with more than 5 times the amount of data. PLR not only emphasizes the necessity of sufficient evidence but also takes into account time perception, visual content perception, and logical reasoning, aiming to build a more advanced video reasoning paradigm.
@@ 2.1 Multimodal Chain-Of-Thought
Multimodal Chain-of-Thought (MCoT) simulates human reasoning by decomposing complex problems into intermediate steps, thereby enhancing transparency and performance. LLaVA-CoT [xu2024llava] is the first to implement a multimodal O1-like [openai2024o1] reasoning approach. Similarly, Video-of-Thought [feivideo] proposes a comprehensive five-stage framework. VideoEspresso [han2025videoespresso] and VideoCoT [wang2024videocot] achieve MCoT by invoking external tools as a multi-stage approach. Chain-of-Frames [ghazanfari2025chain] further integrates this capability into the model’s internal thinking process, enabling it to describe keyframe content. However, single-frame descriptions are static, making it difficult to perform dynamic analysis across frames. We argue that the video modality necessitates CoT because videos contain vast amounts of information, requiring the model to identify crucial details while filtering out noise. The model’s process of progressively identifying key information is inherently dynamic, and our designed perception and analysis loop supports this process.
@@ 2.2 Reasoning VideoLLMs
VideoR1 [feng2025video] first achieves R1-like [guo2025deepseek] video reasoning and utilizes GRPO [shao2024deepseekmath] for training. However, its cold-start data is sampled from Qwen2.5VL 72B [bai2025qwen25vltechnicalreport]. This strategy overlooks the fact that even models with larger parameter counts lack video reasoning training. To enhance video perception, VideoR1 and DeepVideo-R1 [park2025deepvideo] enhance the model’s perception by modifying the standard GRPO. VideoRFT [wang2025videorft] advances this further by restructuring the CoT into a fixed pipeline: task, perception, reasoning, and conclusion. It utilizes SigLip [tschannen2025siglip] during the perception stage to mitigate hallucinations. TimeR1 [wang2025time] and VideoChatR1 [li2025videochat] place a greater emphasis on temporal perception and often lack consideration for general-purpose tasks. VideoChatR1.5 [yan2025videochat] achieves tool-use for video trimming based on temporal reasoning. ARC-HunyuanVideo [ge2025archunyuanvideo7bstructuredvideocomprehension] can generate dense video descriptions with timestamps, but lacks the ability to reason within these segments, and also does not address the hallucination problem. In contrast, our reasoning paradigm integrates temporal perception and execution, anti-hallucination training, and cyclical evidence gathering, representing a more advanced approach to reasoning training.
@@ 2.3 Video Hallucination
Longer reasoning chains reduce attention to visual input, thereby leading to hallucinations [liu2025more]. Consequently, hallucination control is crucial for video models. VideoRFT [wang2025videorft] utilizes SigLip [tschannen2025siglip]. Nevertheless, SigLip is not optimally suited for video. Utilizing a VideoLLM as a judge requires the model itself to be low-hallucinating and unbiased. However, the current models cannot achieve this. VidHalluc [li2025vidhalluc] evaluates hallucinations from three key dimensions, while VideoHallucer [wang2024videohallucer] focuses on the bias of judging the correctness of descriptions. To address this, we provide a controllable data generation pipeline for hallucination pairs and introduce a video hallucination judgement model. Our primary objective is to supply a reliable anti-hallucination reward.
@@ 3 Factual-Aware Evaluator
To provide reliable consistency between textual descriptions and visual content, Factual-Aware Evaluator is trained on AnetHallu-117K, which is based on human-annotated ActivityNet (ANet) [caba2015activitynet]. In this section, we describe in detail the caption construction pipeline and the preference training pipeline, as shown at the top of Fig. 2. The key to this pipeline is to eliminate the bias of language and strengthen the visual judgment of the model.
@@ 3.3 Hybrid Preference Training
Odds Ratio Preference Optimization (ORPO) [hong2024orpo] is employed for training. The total loss function $\mathcal{L}_{ORPO}$ for ORPO is defined as a weighted sum of the SFT loss and the Odds Ratio (OR) preference loss:
@@ 4.2 Cold-Start Data Construction
Video perception is divided into sequential and disordered ones on the time axis. To accommodate the reasoning for these two sequences, we have covered two types of tasks. For the sequential part, we select NextQA [xiao2021next], which contains causal and temporal action reasoning. For the disordered part, we set event ordering based on ActivityNet [caba2015activitynet] named ReOrdering. The challenge is the lack of timestamp and reasoning annotations within NextQA. Thus, we mainly introduce how to construct the CoT with timestamps of NextQA, as shown at the bottom of Fig. 2(b).
@@ 4.3 Training with GRPO and Reward Design
The Group Relative Policy Optimization (GRPO) [shao2024deepseekmath] process begins by sampling $K$ candidate responses $\{o_{i}\}_{i=1}^{K}$ for a given query $q$ and evaluating their rewards $\{r_{i}\}_{i=1}^{K}$. The rewards are then normalized to compute the advantage $A_{i}$ for each response:
@@
Accuracy Reward: Our training data mainly contains five kinds of tasks: Video Temporal Grounding (VTG), Multiple Choice (MC), MC&VTG (GLUE), ReOrdering (RO), and Open-Ended (OE), illustrated in Fig. 3. Except for the OE task, we use rule matching. And for the OE task, we use an open-source verification model (VM) [xVerify]. Given the question $Q$, the ground truth answer $G$, and the model’s generated answer $A_{\text{gen}}$ as input, we obtain the token probabilities $P_{Ic}$ and $P_{C}$ for the Incorrect and Correct labels, respectively. Thus, the accuracy reward $R_{acc}$ is following:
@@
GPT-4o [openai2024gpt4ocard] - 42.0 67.4 46.9 34.0 46.6 71.9 - -
Gemini-2.0-Flash [comanici2025gemini] - 30.6 65.9 51.7 45.4 - - - -
@@
Qwen2.5VL [baiQwen25VLTechnicalReport2025] - 17.0 53.0 24.8 28.6 36.9 52.1 14.3 32.4
Time-R1 [wang2025time] 0+2.5K 32.2 52.5 22.9 28.5 34.9 53.4 21.0 35.3
Video-RFT [wang2025videorft] 102+8K 34.3 52.6 25.8 31.9 35.2 54.7 10.6 34.7
@@
Qwen2.5VL [baiQwen25VLTechnicalReport2025] - 27.8 63.3 29.7 31.8 39.9 60.1 18.7 38.8
Time-R1 [wang2025time] 0+2.5K 37.1 64.2 29.6 28.4 39.0 60.9 29.6 41.3
Video-R1 [feng2025video] 165+8K 36.5 64.2 31.3 35.8 39.1 59.3 13.7 40.0
Video-RFT [wang2025videorft] 102+8K 39.8 68.5 30.6 36.8 40.3 59.8 15.4 41.6
Videochat-R1 [li2025videochat] 0+18K 33.0 63.8 30.0 31.3 39.1 60.8 30.1 41.2
Videochat-R1.5 [yan2025videochat] 0+80K 38.5 64.5 30.0 40.6 41.3 61.2 26.8 43.2
@@
Model VideoHallucer [wang2024videohallucer] HEAVEN [gao2025exploringhallucinationlargemultimodal] VidHalluc  [li2025vidhalluc]
@@
GPT-4o [openai2024gpt4ocard] 74.6 0.9 - - 91.0
Gemini-1.5-Pro [geminiteam2024gemini15unlockingmultimodal] 62.9 41.3 - - 79.2
Qwen2.5VL [baiQwen25VLTechnicalReport2025] 67.8 22.4 63.7 28.9 89.2
@@ 5.1 Setups
Training Details For the Factual-Aware Evaluator, we use LoRA [hu2022lora] based on Qwen2.5VL 7B, training for one epoch with a learning rate of 1e-4, where $\lambda$ in formula 4.3 is 0.5, with a maximum of 16 frames and the total number of pixels limited to $1200\texttimes 28\texttimes 28$. For the reasoning model, training includes a cold-start SFT and RL using GRPO. The cold-start data incorporates 20K image CoT data from the Reason-RFT-CoT-Dataset [tan2025reason] and our VideoPLR-14K. We collected 6K questions for training, sourced from LLava-Video-178K [llava-video], Time-R1 [wang2025time], Cinepile [rawal2024cinepilelongvideoquestion], NextGQA [xiao2024can], and ANet [caba2015activitynet], as show in Fig. 3. Both SFT and RL, the learning rate is set to 2e-6, and train for one epoch. We set the number of generations to 8 and the batch size to 6, the $\beta$ eq. 4.3 is set to 0.01, and the maximum output length is set to 1600 tokens. The maximum number of frames and total pixels are set to $64$ and $2048\times 28\times 28$. Our training code is optimized based on ms-swift [zhao2024swiftascalablelightweightinfrastructure] to adapt to the anti-hallucination reward, which incurs significant computational latency. Thus, we implement GRPO training without additional time consumption.
Evaluation Details First, we evaluate the Factual-Aware Evaluator, including the VidHalluc [li2025vidhalluc], VideoHallucer [wang2024videohallucer], and HEAVEN [gao2025exploringhallucinationlargemultimodal]. Set the maximum number of frames to 16, the number of pixels to $2400\times 28\times 28$, and the temperature to 0. For video reasoning models, the evaluation benchmarks include three types: reasoning, general, and temporal. The reasoning benchmarks include: Video-Holmes [cheng2025videoholmesmllmthinklike], MMVU [zhao2025mmvu], VCR-benchmark [qi2025vcrbenchcomprehensiveevaluationframework],Video-tt [zhang2025towards], VSI-bench [yang2025thinking]. The general benchmark is VideoMME [fu2025video], and the temporal benchmark is TVG-Bench [wang2025time]. Set the maximum number of frames to 64, the number of pixels to $6000\times 28\times 28$, and the temperature to 0.
@@ 5.2 Hallucination Evaluation
The evaluation begins with the hallucination benchmark evaluation, as shown in Tab 2. We use VideoHallucer [wang2024videohallucer] for binary judgment, the open-source HEAVEN [gao2025exploringhallucinationlargemultimodal] data reformatted as captions, and the MCQ task from VidHalluc [li2025vidhalluc]. On VideoHallucer, FAE surpasses GPT-4o. It boosts accuracy over the base model by 7.5% and 9.2%. FAE improves judgment fairness, reducing bias towards ‘Y’ or ‘N’. It narrows the accuracy gap between correct and incorrect classifications by 15.1% and 26.8%. This reduction validates our data construction method and FAE’s reliability as an anti-hallucination reward, confirming the robustness and successful transfer of its learned abilities. Despite being trained exclusively on binary classification tasks, our model demonstrates strong generalization to the MCQ task. This indicates that the learned ability to judge hallucinations is robust and successfully transferred, validating the effectiveness of our training approach.

### v2

@@ Abstract
Sufficient visual perception is the foundation of video reasoning. Nevertheless, existing Video Reasoning LLMs suffer from perception shortcuts, relying on a flawed single-step perception paradigm. This paradigm describes the video and then conducts reasoning, which runs the risk of insufficient evidence and emergent hallucinations. To address these issues, we introduce a new framework that integrates a loop-based paradigm with an anti-hallucination reward. First, to address the insufficient evidence, we introduce the Perception Loop Reasoning (PLR) paradigm. Instead of describing the video at once, each loop requires the model to describe a video segment with precise timestamps, analyze this segment, and decide the next action. Second, for the risk of hallucinations, the Factual-Aware Evaluator (FAE) evaluates each perception result as a reliable anti-hallucination reward. This reward encourages the model to provide sufficient and precise video evidence. Our FAE, which performs comparably to GPT-4o, is tuned on our AnetHallu-117K, a large-scale hallucination judgment preference dataset. Extensive experiments show that our Video-PLR achieves the state-of-the-art in both 3B and 7B parameter scales and has the best data efficiency. Our code, datasets, and models are released on: https://github.com/BoweiPu/VideoPLR.
@@ 1 Introduction
Video Large Language Models (VideoLLMs) [45, 21, 33] demonstrate remarkable potential in traditional video understanding. Building on this success, researchers are increasingly turning their attention to complex video tasks. Inspired by the Think-then-Answer paradigm [17], many VideoLLMs [32, 43, 12] tackle complex problems by generating intermediate reasoning steps. This reasoning is based on video perception: the accurate description of visual information in the video. Thus, insufficient perception or hallucinations often lead to flawed or erroneous reasoning.
@@
Realizing this multi-step perception is challenging, as current models lack the ability to re-perceive, and no dataset exists to support it. Following the PLR paradigm, we construct VideoPLR-14K, a novel dataset designed to bootstrap the model to generate responses in the PLR format, encompassing both sequential causal reasoning and non-sequential segment ordering. Furthermore, anti-hallucination is the key to precise perception, which is the foundation of reasoning. However, previous models, including Open-source VideoLLMs [34, 22] and vision-language models [42], cannot provide reliable factual evaluation. To fill this gap, we introduce a Factual-Aware Evaluator (FAE), which is meticulously trained on our newly constructed AnetHallu-117K dataset. The ANetHallu-117K is constructed from the human-annotated ANet [10] and contains 5 types of hallucination and hybrid CoT binary classification preference data, which can reduce discriminative bias. FAE is comparable to GPT-4o in evaluating text-video consistency to serve as a reliable anti-hallucination reward for RL training. The segmented approach of PLR is also more reliable for hallucination evaluation. Specifically, their constrained visual context simplifies grounding the descriptions, making anti-hallucination more readily identifiable.
These two efforts jointly constitute our final model, Video-PLR via reinforcement learning [17], achieving high data efficiency. For the RL stage, we utilize only 6K of mixed data, including five types of tasks. Our models (3B and 7B) achieve SOTA results on multiple benchmarks, as illustrated in Fig. 1(c). Compared with previous models, Video-PLR integrates accurate timestamps directly into the reasoning text, enabling a unified consideration of temporal grounding. It is worth mentioning that by simply adding 0.5K of Video Temporal Grounding data, we outperform models trained with more than 5 times the amount of data. PLR not only emphasizes the necessity of sufficient evidence but also takes into account time perception, visual content perception, and logical reasoning, aiming to build a more advanced video reasoning paradigm.
@@ 2.1 Multimodal Chain-Of-Thought
Multimodal Chain-of-Thought (MCoT) simulates human reasoning by decomposing complex problems into intermediate steps, thereby enhancing transparency and performance. LLaVA-CoT [39] is the first to implement a multimodal O1-like [25] reasoning approach. Similarly, Video-of-Thought [11] proposes a comprehensive five-stage framework. VideoEspresso [18] and VideoCoT [35] achieve MCoT by invoking external tools as a multi-stage approach. Chain-of-Frames [16] further integrates this capability into the model’s internal thinking process, enabling it to describe keyframe content. However, single-frame descriptions are static, making it difficult to perform dynamic analysis across frames. We argue that the video modality necessitates CoT because videos contain vast amounts of information, requiring the model to identify crucial details while filtering out noise. The model’s process of progressively identifying key information is inherently dynamic, and our designed perception and analysis loop supports this process.
@@ 2.2 Reasoning VideoLLMs
VideoR1 [12] first achieves R1-like [17] video reasoning and utilizes GRPO [29] for training. However, its cold-start data is sampled from Qwen2.5VL 72B [1]. This strategy overlooks the fact that even models with larger parameter counts lack video reasoning training. To enhance video perception, VideoR1 and DeepVideo-R1 [26] enhance the model’s perception by modifying the standard GRPO. VideoRFT [32] advances this further by restructuring the CoT into a fixed pipeline: task, perception, reasoning, and conclusion. It utilizes SigLip [31] during the perception stage to mitigate hallucinations. TimeR1 [36] and VideoChatR1 [23] place a greater emphasis on temporal perception and often lack consideration for general-purpose tasks. VideoChatR1.5 [40] achieves tool-use for video trimming based on temporal reasoning. ARC-HunyuanVideo [15] can generate dense video descriptions with timestamps, but lacks the ability to reason within these segments, and also does not address the hallucination problem. In contrast, our reasoning paradigm integrates temporal perception and execution, anti-hallucination training, and cyclical evidence gathering, representing a more advanced approach to reasoning training.
@@ 2.3 Video Hallucination
Longer reasoning chains reduce attention to visual input, thereby leading to hallucinations [24]. Consequently, hallucination control is crucial for video models. VideoRFT [32] utilizes SigLip [31]. Nevertheless, SigLip is not optimally suited for video. Utilizing a VideoLLM as a judge requires the model itself to be low-hallucinating and unbiased. However, the current models cannot achieve this. VidHalluc [22] evaluates hallucinations from three key dimensions, while VideoHallucer [34] focuses on the bias of judging the correctness of descriptions. To address this, we provide a controllable data generation pipeline for hallucination pairs and introduce a video hallucination judgement model. Our primary objective is to supply a reliable anti-hallucination reward.
@@ 3 Factual-Aware Evaluator
To provide reliable consistency between textual descriptions and visual content, Factual-Aware Evaluator is trained on AnetHallu-117K, which is based on human-annotated ActivityNet (ANet) [10]. In this section, we describe in detail the caption construction pipeline and the preference training pipeline, as shown at the top of Fig. 2. The key to this pipeline is to eliminate the bias of language and strengthen the visual judgment of the model.
@@ 3.3 Hybrid Preference Training
Odds Ratio Preference Optimization (ORPO) [19] is employed for training. The total loss function $\mathcal{L}_{ORPO}$ for ORPO is defined as a weighted sum of the SFT loss and the Odds Ratio (OR) preference loss:
@@ 4.2 Cold-Start Data Construction
Video perception is divided into sequential and disordered ones on the time axis. To accommodate the reasoning for these two sequences, we have covered two types of tasks. For the sequential part, we select NextQA [37], which contains causal and temporal action reasoning. For the disordered part, we set event ordering based on ActivityNet [10] named ReOrdering. The challenge is the lack of timestamp and reasoning annotations within NextQA. Thus, we mainly introduce how to construct the CoT with timestamps of NextQA, as shown at the bottom of Fig. 2(b).
@@ 4.3 Training with GRPO and Reward Design
The Group Relative Policy Optimization (GRPO) [29] process begins by sampling $K$ candidate responses $\{o_{i}\}_{i=1}^{K}$ for a given query $q$ and evaluating their rewards $\{r_{i}\}_{i=1}^{K}$. The rewards are then normalized to compute the advantage $A_{i}$ for each response:
@@
Accuracy Reward: Our training data mainly contains five kinds of tasks: Video Temporal Grounding (VTG), Multiple Choice (MC), MC&VTG (GLUE), ReOrdering (RO), and Open-Ended (OE), illustrated in Fig. 3. Except for the OE task, we use rule matching. And for the OE task, we use an open-source verification model (VM) [3]. Given the question $Q$, the ground truth answer $G$, and the model’s generated answer $A_{\text{gen}}$ as input, we obtain the token probabilities $P_{Ic}$ and $P_{C}$ for the Incorrect and Correct labels, respectively. Thus, the accuracy reward $R_{acc}$ is following:
@@
GPT-4o [9] - 42.0 67.4 46.9 34.0 46.6 71.9 - -
Gemini-2.0-Flash [7] - 30.6 65.9 51.7 45.4 - - - -
@@
Qwen2.5VL [2] - 17.0 53.0 24.8 28.6 36.9 52.1 14.3 32.4
Time-R1 [36] 0+2.5K 32.2 52.5 22.9 28.5 34.9 53.4 21.0 35.3
Video-RFT [32] 102+8K 34.3 52.6 25.8 31.9 35.2 54.7 10.6 34.7
@@
Qwen2.5VL [2] - 27.8 63.3 29.7 31.8 39.9 60.1 18.7 38.8
Time-R1 [36] 0+2.5K 37.1 64.2 29.6 28.4 39.0 60.9 29.6 41.3
Video-R1 [12] 165+8K 36.5 64.2 31.3 35.8 39.1 59.3 13.7 40.0
Video-RFT [32] 102+8K 39.8 68.5 30.6 36.8 40.3 59.8 15.4 41.6
Videochat-R1 [23] 0+18K 33.0 63.8 30.0 31.3 39.1 60.8 30.1 41.2
Videochat-R1.5 [40] 0+80K 38.5 64.5 30.0 40.6 41.3 61.2 26.8 43.2
@@
Model VideoHallucer [34] HEAVEN [14] VidHalluc  [22]
@@
GPT-4o [9] 74.6 0.9 - - 91.0
Gemini-1.5-Pro [8] 62.9 41.3 - - 79.2
Qwen2.5VL [2] 67.8 22.4 63.7 28.9 89.2
@@ 5.1 Setups
Training Details For the Factual-Aware Evaluator, we use LoRA [20] based on Qwen2.5VL 7B, training for one epoch with a learning rate of 1e-4, where $\lambda$ in formula 4.3 is 0.5, with a maximum of 16 frames and the total number of pixels limited to $1200\texttimes 28\texttimes 28$. For the reasoning model, training includes a cold-start SFT and RL using GRPO. The cold-start data incorporates 20K image CoT data from the Reason-RFT-CoT-Dataset [30] and our VideoPLR-14K. We collected 6K questions for training, sourced from LLava-Video-178K [45], Time-R1 [36], Cinepile [28], NextGQA [38], and ANet [10], as show in Fig. 3. Both SFT and RL, the learning rate is set to 2e-6, and train for one epoch. We set the number of generations to 8 and the batch size to 6, the $\beta$ eq. 4.3 is set to 0.01, and the maximum output length is set to 1600 tokens. The maximum number of frames and total pixels are set to $64$ and $2048\times 28\times 28$. Our training code is optimized based on ms-swift [47] to adapt to the anti-hallucination reward, which incurs significant computational latency. Thus, we implement GRPO training without additional time consumption.
Evaluation Details First, we evaluate the Factual-Aware Evaluator, including the VidHalluc [22], VideoHallucer [34], and HEAVEN [14]. Set the maximum number of frames to 16, the number of pixels to $2400\times 28\times 28$, and the temperature to 0. For video reasoning models, the evaluation benchmarks include three types: reasoning, general, and temporal. The reasoning benchmarks include: Video-Holmes [6], MMVU [48], VCR-benchmark [27],Video-tt [46], VSI-bench [41]. The general benchmark is VideoMME [13], and the temporal benchmark is TVG-Bench [36]. Set the maximum number of frames to 64, the number of pixels to $6000\times 28\times 28$, and the temperature to 0.
@@ 5.2 Hallucination Evaluation
The evaluation begins with the hallucination benchmark evaluation, as shown in Tab 2. We use VideoHallucer [34] for binary judgment, the open-source HEAVEN [14] data reformatted as captions, and the MCQ task from VidHalluc [22]. On VideoHallucer, FAE surpasses GPT-4o. It boosts accuracy over the base model by 7.5% and 9.2%. FAE improves judgment fairness, reducing bias towards ‘Y’ or ‘N’. It narrows the accuracy gap between correct and incorrect classifications by 15.1% and 26.8%. This reduction validates our data construction method and FAE’s reliability as an anti-hallucination reward, confirming the robustness and successful transfer of its learned abilities. Despite being trained exclusively on binary classification tasks, our model demonstrates strong generalization to the MCQ task. This indicates that the learned ability to judge hallucinations is robust and successfully transferred, validating the effectiveness of our training approach.